A 200 g calorimeter can contains 150 g oil at temperature of 20^0 C. To the oil is added 80 g of aluminum at 300^0C, what will be the temperature of the system after equilibrium is established? Ccu=0.093 cal/g^0C
Coil=0.37 cal/g^0C
CAl=0.21 cal/g^0C
1
Expert's answer
2018-01-31T05:26:07-0500
For the thermal equilibrium:
m_ca C_ca (T-T_1)+m_oil C_oil (T-T_1)=m_Al C_Al (T_2-T)
(200)(0.093)(T-20)+(150)(0.37)(T-20)=(80)(0.21)(300-T)
The temperature of the system after equilibrium is established is
T=72℃.